emmagrace231 (emmagrace231):

Can someone please help me with this question? I don't know the formula or how to find it.

emmagrace231 (emmagrace231):

this is the question. I have to do number 4 only.

OpenStudy (texaschic101):

(-2,3).....x = -2 and y = 3 slope(m) = -2 y = mx + b 3 = -2(-2) + b 3 = 4 + b 3 - 4 = b -1 = b so your equation is : y = -2x + (-1)...or y = -2x - 1

OpenStudy (texaschic101):

oh..wait....point slope...hold on

OpenStudy (texaschic101):

y - y1 = m(x - x1) (-2,3)....x1 = -2 and y1 = 3 slope(m) = -2 now we sub y - 3 = -2(x - (-2) = y - 3 = -2(x + 2)

OpenStudy (texaschic101):

point slope formula is : y - y1 = m(x - x1)